# Gueter Josmy Faure1    Min-Hung Chen2    Jia-Fong Yeh1    Hung-Ting Su1 Shang-Hong Lai4    Winston H. Hsu1,3

### v3

@@ A.2 Implementation Details
A.3 Model Details
A.3.1 Details of our Episodic QFormer
@@
A.3.2 Details of SeTR
@@ A.4 Extended Ablations
A.4.1 How does the number of frames affect the model’s accuracy and latency?
MovieChat (Song et al., 2024) processes 2048 frames for each video, while we use only 100 frames, as previous studies have demonstrated how redundant video data is (Simonyan & Zisserman, 2014; Wang et al., 2016). Given that the MovieChat-1k dataset contains very long videos (some exceeding 14,000 frames), we conducted experiments to extend the number of frames our model processes. Specifically, we experiment with 40, 80, 100, 300, 500, and 1000 frames while keeping the number of episodes constant. As for the SeTR keep ratio, we decrease it in function of the number frames so that the number of semantic features we keep equals 20.
@@ A.4.1 How does the number of frames affect the model’s accuracy and latency?
A.4.2 How does the window size affect the model’s accuracy and latency?
@@
A.4.3 A Note on Latency
@@
A.5 More Qualitative Results
To further illustrate the capabilities of our model, we present a series of qualitative examples that highlight its strengths in various long-form video understanding tasks.
@@ A.5 More Qualitative Results
A.6 Error Analysis: When does HERMES fail and why?
@@
A.7 How is our approach related to cognitive processes?
Our approach to long-form video understanding is inspired by cognitive processes involving memory and comprehension. According to the literature on neuroscience (Tulving et al., 1972; Schacter & Tulving, 1982; Tulving, 1983), human cognition involves two primary types of memory: episodic and semantic. Episodic memory is the ability to recall specific events or episodes, while semantic memory refers to the storage of general knowledge and concepts. These forms of memory are crucial for understanding long-form narratives, where a coherent understanding arises from the integration of specific events and overarching themes.
The proposed HERMES model incorporates these cognitive processes through its two main components, ECO and SeTR. ECO, akin to the function of episodic memory, selectively retains and compresses key events from the video, allowing the model to form a structured representation of the narrative as it unfolds. This approach is an oversimplified abstraction of findings in cognitive neuroscience, which highlight the role of the hippocampus in the consolidation of episodic memories (Eichenbaum, 2004; Schacter & Tulving, 1982), and the concept of subjective time (Arstila et al., 2014) that sees a scene (or a video) not as a series of frames but as a series of experiences. The hippocampus enables the organization of temporally distinct experiences into a coherent memory trace, something that we aim to capture with ECO. Moreover, the sequential processing and aggregation of information in our model align with the concept of event segmentation in cognitive psychology (Zacks et al., 2007). Humans naturally segment continuous experiences into discrete events, which aids in memory formation and recall.
Meanwhile, SeTR functions similarly to semantic memory, extracting and reinforcing high-level semantic cues. This process mirrors how the brain integrates detailed episodic memories with broader semantic knowledge stored in the neocortex (McClelland et al., 1995; Binder & Desai, 2011). Also related is the concept of gist extraction which involves rapidly comprehending the essence or overall meaning of a scene or situation (Oliva, 2005). This ability allows humans to quickly understand the context of a complex scene without processing every detail. Our SeTR operates similarly by identifying and extracting high-level semantic cues that provide a concise overview of the scene and actions.

### v4

@@ H.2 Implementation Details
H.3 Model Details
H.3.1 Details of our Episodic QFormer
@@
To mitigate temporal confusion during merging, we apply positional encoding (PE) to frame features before ECO. This effectively discourages out-of-order merges by embedding temporal locality directly into similarity calculations. As an ablation, removing PE reduces MovieChat-1k accuracy from 78.6 to 77.3 on MovieChat-1k, indicating its effectiveness in preserving temporal coherence despite compression. Other studies such as Transformer-XL [8] and Compressive Transformer [28], also report performance drops when positional biases are removed from their compression modules.
ECO implicitly captures event frequency: frequent events naturally occur across multiple frames and thus have higher likelihoods of being retained or merged into reinforced prototypes within the memory bank. This self-reinforcing mechanism ensures high-importance (and often high-frequency) events remain well-represented. Explicit event frequency tracking is an idea worth exploring, however, we believe it would be more computationally intensive and may force important but infrequent representations out of memory.
H.3.2 Details of SeTR
@@ H.4 Extended Ablations
H.4.1 How does the number of frames affect the model’s accuracy and latency?
MovieChat [32] processes 2048 frames for each video, while we use only 100 frames, as previous studies have demonstrated how redundant video data is [31, 42]. Given that the MovieChat-1k dataset contains very long videos (some exceeding 14,000 frames), we conducted experiments to extend the number of frames our model processes. Specifically, we experiment with 40, 80, 100, 300, 500, and 1000 frames while keeping the number of episodes constant. As for the SeTR keep ratio, we decrease it in function of the number frames so that the number of semantic features we keep equals 20.
@@ H.4.1 How does the number of frames affect the model’s accuracy and latency?
H.4.2 How does the window size affect the model’s accuracy and latency?
@@
H.5 HERMES vs. MA-LMM vs. MovieChat
HERMES versus MA-LMM: For each incoming frame, MA-LMM adds it to the memory bank by computing the similarities with adjacent frames and merging the incoming frame with its most similar in the memory bank. Below are our main differences.
  • HERMES takes a distributed approach. Our ECO, distributes the frames of the incoming window to the most appropriate episode. This approach is more intuitive and better mirrors human memory formation.
  • Frames can be grouped into episodes regardless of temporal adjacency, unlike MA-LMM which only considers adjacent frames. This naturally handles scene transitions, flashbacks, and non-linear narratives.
  • HERMES is vastly more efficient and accurate. As shown in Table 5 in the main paper, our memory management system almost halves the inference time (-43%) when plugged into MA-LMM while being 3.4% more accurate.
  • HERMES  also captures semantics. Our Semantics Retriever (SeTR) complements the episodic memory and is shown in Table 5 to increase the accuracy of MA-LMM by almost 4% with only a negligible increase in latency.
HERMES versus MovieChat: Moviechat’s short-term memory uses a FIFO mechanism. Its long-term memory uses ToMe. Below are the main differences
[…]
  • HERMES’s long-term memory is implicitly encoded in ECO. We consider SeTR as a semantics scanner that retrieves scattered semantics from the video.
  • 22 FPS processing speed compared to MovieChat’s  0.01 FPS (13 minutes vs 1 day on MovieChat-1k) using a V100 GPU (32 GB).
  • HERMES  achieves high performance with only 100 frames compared to MovieChat’s 2048 frames.
H.6 A Note on Latency
@@
H.7 Qualitative Results
@@
H.7.1 Visualization of ECO and SeTR
Figure 12 demonstrates the inner-workings of ECO and SeTR. The top row illustrates a curated summary of the video content, highlighting diverse scenes, such as landscapes, wildlife, and environmental features.
SeTR is responsible for extracting high-level semantic features and grouping frames with similar themes, as shown in the mid row. For instance, the module effectively captures thematic clusters such as “Landscape,” “Various Birds,” and “Reptiles,” providing a concise overview of the video.
Meanwhile, ECO processes the video at a more granular level, segmenting it into coherent episodes that reflect the narrative flow. The bottom row showcases this segmentation, organizing the content into episodic units like “Arid Landscape,” “Lake and Aquatic Bird,” and “Flies.” This two-tiered approach ensures both thematic abstraction and temporal coherence, enabling a comprehensive understanding of the video.
Figure 12: Visualization of ECO and SeTR: The top row presents a curated visual summary of the video, showcasing key scenes such as landscapes, wildlife, and environmental features. The middle row highlights the functionality of SeTR, which extracts semantic features and clusters frames into thematic groups, including “Landscape,” “Various Birds,” and “Reptiles.” Finally, the bottom row illustrates the operation of ECO, which segments the video into coherent narrative episodes, such as “Arid Landscape,” “Lake and Aquatic Bird,” and “Flies.” Together, these modules provide both high-level abstraction and detailed episodic structure for comprehensive video understanding.
H.8 Error Analysis: When does HERMES fail and why?
@@
H.9 How is our approach related to cognitive processes?
Our approach to long-form video understanding is inspired by cognitive processes involving memory and comprehension. According to the literature on neuroscience [39, 30, 38], human cognition involves two primary types of memory: episodic and semantic. Episodic memory is the ability to recall specific events or episodes, while semantic memory refers to the storage of general knowledge and concepts. These forms of memory are crucial for understanding long-form narratives, where a coherent understanding arises from the integration of specific events and overarching themes.
The proposed HERMES model incorporates these cognitive processes through its two main components, ECO and SeTR. ECO, akin to the function of episodic memory, selectively retains and compresses key events from the video, allowing the model to form a structured representation of the narrative as it unfolds. This approach is an oversimplified abstraction of findings in cognitive neuroscience, which highlight the role of the hippocampus in the consolidation of episodic memories [9, 30], and the concept of subjective time [1] that sees a scene (or a video) not as a series of frames but as a series of experiences. The hippocampus enables the organization of temporally distinct experiences into a coherent memory trace, something that we aim to capture with ECO. Moreover, the sequential processing and aggregation of information in our model align with the concept of event segmentation in cognitive psychology [47]. Humans naturally segment continuous experiences into discrete events, which aids in memory formation and recall.
Meanwhile, SeTR functions similarly to semantic memory, extracting and reinforcing high-level semantic cues. This process mirrors how the brain integrates detailed episodic memories with broader semantic knowledge stored in the neocortex [23, 2]. Also related is the concept of gist extraction which involves rapidly comprehending the essence or overall meaning of a scene or situation [26]. This ability allows humans to quickly understand the context of a complex scene without processing every detail. Our SeTR operates similarly by identifying and extracting high-level semantic cues that provide a concise overview of the scene and actions.